OpenStudy (anonymous):

can someone check my answers for a couple problems? Please(: Find the vertices and foci of the hyperbola with equation quantity ((x+4)^2/9) - ((y+3)^2/16)=1 * I got vertices:(-1,-3),(-7,-3) focie:(-9,-3)(1,-3)

OpenStudy (anonymous):

another one is, Find the center, vertices, and foci of the ellipse with equation 3x2 + 8y2 = 24. * i got Center: (0, 0); Vertices: (-2sqrt2,0)(2sqrt2,0) foci:(-sqrt5,0)(sqrt5,0)

OpenStudy (anonymous):

and my last one, Find the center, vertices, and foci of the ellipse with equation (x^2)/9+(y^2)/25 = 1. * my answer was center:(0,0) vertices:(-5,0)(5,0) foci:(-4,0)(4,0)

OpenStudy (jdoe0001):

your values for the 1st and 2nd are good the value for the last one, are good too, they just need a minor tweak \(\bf \cfrac{x^2}{3^2}+\cfrac{y^2}{5^2}=1\) in an ellipse, the major traverse axis is where the bigger denominator is in this case, the fraction with the y-component so you'd just need to swap your values from \(\bf vertices:(-5,0)(5,0) foci:(-4,0)(4,0)\) to \(\bf vertices:(0, -5)(0, 5) foci:(0,-4)(0, 4)\)

OpenStudy (anonymous):

thankyou!(:

OpenStudy (jdoe0001):

yw